OpenStudy (anonymous):

what is the value of the expression? 6 + 3 x 2 / 3 x 4 - 10

OpenStudy (ageta):

Let's simplify step-by-step. 6+ 3x(2) 3 x(4)−10 =6+8x2+−10 Combine Like Terms: =6+8x2+−10 =(8x2)+(6+−10) =8x2+−4 Answer: =8x2−4